# Inferring Change Points in High-Dimensional Regression via Approximate Message Passing

### v2

@@ Abstract
We consider the problem of localizing change points in a generalized linear model (GLM), a model that covers many widely studied problems in statistical learning including linear, logistic, and rectified linear regression. We propose a novel and computationally efficient Approximate Message Passing (AMP) algorithm for estimating both the signals and the change point locations, and rigorously characterize its performance in the high-dimensional limit where the number of parameters $p$ is proportional to the number of samples $n$. This characterization is in terms of a state evolution recursion, which allows us to precisely compute performance measures such as the asymptotic Hausdorff error of our change point estimates, and allows us to tailor the algorithm to take advantage of any prior structural information on the signals and change points. Moreover, we show how our AMP iterates can be used to efficiently compute a Bayesian posterior distribution over the change point locations in the high-dimensional limit. We validate our theory via numerical experiments, and demonstrate the favorable performance of our estimators on both synthetic and real data in the settings of linear, logistic, and rectified linear regression.
Keywords: change point localization, high-dimensional regression, generalized linear models, multi-index models, approximate message passing
@@ 1 Introduction
Heterogeneity is a common feature of large, high-dimensional datasets. When the data are ordered by time, a simple form of heterogeneity is a change in the data generating mechanism at certain unknown instants of time. If these ‘change points’ were known, or estimated accurately, the dataset could be partitioned into homogeneous subsets, each amenable to analysis via standard statistical techniques (Fryzlewicz, 2014). Models with change points have been studied in a variety of statistical contexts, such as the detection of changes in: signal means (Wang et al., 2020; Liu et al., 2021; Li et al., 2024); covariance structures (Cho and Fryzlewicz, 2015; Wang et al., 2021b); graphs (Londschien et al., 2023; Bhattacharjee et al., 2020; Fan and Guan, 2018); dynamic networks (Wang et al., 2021a); and functionals (Madrid Padilla et al., 2022). Change point models have found application in a range of fields including genomics (Braun et al., 2000), neuroscience (Aston and Kirch, 2012), causality (He et al., 2022), and economics (Andreou and Ghysels, 2002).
@@
Here, ${\bm{\beta}}^{(i)}\in{\mathbb{R}}^{p}$ is the unknown regression vector (signal) for the $i$th sample, $\bm{X}_{i}\in{\mathbb{R}}^{p}$ is the (known) covariate vector, $\varepsilon_{i}$ is noise, and $q:{\mathbb{R}}\times{\mathbb{R}}\to{\mathbb{R}}$ is a known function. We denote the unknown change points, i.e., the sample indices where the regression vector changes, by $\eta_{1},\ldots,\eta_{L^{*}-1}$. Specifically, we have
$1=\eta_{0}<\eta_{1}<\dots<\eta_{L^{*}}=n,$
with ${\bm{\beta}}^{(i)}\neq{\bm{\beta}}^{(i-1)}$ if and only if $i\in\{\eta_{\ell}\}_{\ell=1}^{L^{*}-1}$. We note that $L^{*}$ is the number of distinct signals in the sequence $\{{\bm{\beta}}^{(i)}\}_{i=1}^{n}$, and $(L^{*}-1)$ is the number of change points. The number of change points is not known, but an upper bound $L$ on the value of $L^{*}$ is available. The goal is to estimate the change point locations as well as the $L^{*}$ signals. We would also like to quantify the uncertainty in these estimates, e.g., via confidence sets or a posterior distribution. As we describe below, the model (1) covers many widely studied regression models including linear, logistic, and rectified linear regression.
@@ 1.1 Linear Regression with Change Points
Linear regression with change points in the high-dimensional regime (where the dimension $p$ is comparable to, or exceeds, the number of samples $n$) has been studied in a number of recent works, e.g. Lee et al. (2016); Leonardi and Bühlmann (2016); Kaul et al. (2019); Rinaldo et al. (2021); Xu et al. (2024); Li et al. (2023); Bai and Safikhani (2023). Most of these papers consider the setting where the signals are sparse (the number of non-zero entries in ${\bm{\beta}}^{(i)}\in{\mathbb{R}}^{p}$ is $o(p)$), and analyze procedures that combine the LASSO estimator (or a variant) with a partitioning technique, e.g., dynamic programming. The work of Gao and Wang (2022) assumes sparsity on the difference between signals across a change point, and Cho et al. (2024); Liu et al. (2024) consider general non-sparse signals. The recent works of Liu et al. (2024); Yang et al. (2024) propose change point detection methods for high-dimensional linear models that do not rely on strict sub-Gaussian error assumptions, with performance comparable to that of Xu et al. (2024) in the case of light-tailed additive noise distributions.
@@ 1.2 Logistic Regression with Change Points
Logistic models are widely used for classification, and logistic regression with change points has been studied in number of contexts. In epidemiology, for example, logistic regression with unknown change points has been used to model the relationship between the continuous exposure variable and disease risk (Pastor-Barriuso et al., 2003). In medicine, it has been used to identify relevant immune response biomarkers in patients with potentially infectious diseases (Fong et al., 2015).
@@ 1.3 Rectified Linear Regression with Change Points
Rectified linear regression is an important primitive in the theory of deep learning, and the problem without change points has been extensively studied in machine learning in recent years, both from the perspective of designing new algorithms and from that of computational hardness, e.g.,(Soltanolkotabi, 2017; Diakonikolas et al., 2021).
@@
Although existing procedures for high-dimensional change point regression such as Wang et al. (2023) incorporate sparsity-based constraints, they cannot be easily adapted to take advantage of other kinds of signal priors. Moreover, they are not well-equipped to exploit prior information on the change point locations. Bayesian approaches to change point detection have been studied in several works, e.g. (Fearnhead, 2006; Lungu et al., 2022), however they mainly focus on (low-dimensional) time-series.
@@
Although our theoretical results do not explicitly need assumptions on the separation between adjacent change points, they are most interesting when the separation is of order $n$, i.e., $\Delta:=\min_{\ell\in[L]}(\eta_{\ell}-\eta_{\ell-1})=O(n)$. This separation is natural in our regime, where the number of samples $n$ is proportional to $p$ and the number of degrees of freedom in the signals also grows linearly in $p$. Existing results on high-dimensional change point regression usually assume signals that are $s$-sparse with $s$ sufficiently small, and demonstrate change point estimators that are consistent when $\Delta=\omega\big{(}s\log p/\kappa^{2}\big{)}$, where $\kappa$ is a constant determined by the separation between the signals (Wang et al., 2021c; Li et al., 2023; Wang et al., 2023). In contrast, we do not assume signal sparsity that is sublinear in $n$, so the change point estimation error will not converge to zero unless $n/p\to\infty$. We therefore quantify the AMP performance via precise asymptotics for the estimation error and the limiting posterior distribution.
@@
AMP, a family of iterative algorithms first proposed for linear regression (Kabashima, 2003; Donoho et al., 2009; Krzakala et al., 2012), has been applied to a variety of high-dimensional estimation problems including estimation in generalized linear models (Rangan, 2011; Schniter and Rangan, 2014; Maillard et al., 2020; Mondelli and Venkataramanan, 2021) and low-rank matrix estimation (Fletcher and Rangan, 2018; Lesieur et al., 2017; Montanari and Venkataramanan, 2021; Barbier et al., 2020). An attractive feature of AMP algorithms is that under suitable model assumptions, their performance in the high-dimensional limit can be characterized by a succinct deterministic recursion called state evolution. The state evolution characterization has been used to show that AMP achieves Bayes-optimal performance for some models (Deshpande and Montanari, 2014; Donoho et al., 2013; Barbier et al., 2019).
An important feature of our AMP algorithm, in contrast to the above works, is the use of non-separable denoising functions. (We say a function $g:{\mathbb{R}}^{m\times L}\to{\mathbb{R}}^{m\times L}$ is separable if it acts row-wise identically on the input matrix.) Even with simplifying assumptions on the signal and noise distributions, non-separable AMP denoisers are required to handle the temporal dependence caused by the change points, and allow for precise uncertainty quantification around possible change point locations. Our main state evolution result (Theorem 1) leverages recent results by Berthier et al. (2019) and Gerbelot and Berthier (2023) for AMP with non-separable denoisers. Non-separable denoisers for AMP have been been studied for generalized linear models and for matrix tensor product models, to exploit the dependence within the signal (Som and Schniter, 2012; Metzler et al., 2016; Ma et al., 2019; Rossetti and Reeves, 2023) or between the covariates (Zhang et al., 2024). In our setting, non-separable denoisers are required to exploit the dependence in the observations caused by the change points.
@@ 2 Preliminaries
We let $[n]=\{1,2,\dots,n\}$. We use boldface notation for matrices and vectors. For vectors $\bm{x},\bm{y}\in{\mathbb{R}}^{n}$, we write $\bm{x}\leq\bm{y}$ if $x_{i}\leq y_{i}$ for all $i\in[n]$, and let $[\bm{x},\bm{y}]:=\{\bm{v}\in{\mathbb{R}}^{n}:x_{i}\leq v_{i}\leq y_{i}\;\forall i\in[n]\}$. For a matrix $\bm{A}\in{\mathbb{R}}^{m\times L}$ and $i\in[m],j\in[L]$, we let $\bm{A}_{[i,:]}$ and $\bm{A}_{[:,j]}$ denote its $i$th row and $j$th column respectively. Similarly, for a vector ${\bm{\psi}}\in[L]^{n}$, we let $\bm{A}_{[:,{\bm{\psi}}]}\in{\mathbb{R}}^{m}$ denote the vector whose $i$-th entry is $\bm{A}_{i,\psi_{i}}$. For two sequences (in $n$) of random variables $X_{n}$, $Y_{n}$, we write $X_{n}\stackrel{{\scriptstyle{\mathbb{P}}}}{{\simeq}}Y_{n}$ when their difference converges in probability to 0, i.e., $\lim_{n\to\infty}{\mathbb{P}}(|X_{n}-Y_{n}|>\epsilon)=0$ for any $\epsilon>0$. Denote the covariance matrix of random vector $\bm{Z}\in\mathbb{R}^{q}$ as ${\rm{{Cov}}}(\bm{Z})\in\mathbb{R}^{q\times q}$. We refer to all random elements, including vectors and matrices, as random variables. When referring to probability densities, we include probability mass functions, with integrals interpreted as sums when the distribution is discrete.
@@
Our results are stated in terms of uniformly pseudo-Lipschitz functions (Berthier et al., 2019). For $C>0$ and $r\in[1,\infty)$, let $PL_{n,m,q}(r,C)$ be the set of functions $\phi:{\mathbb{R}}^{n\times q}\to{\mathbb{R}}^{m\times q}$ such that
@@ 3 AMP Algorithm and Main Results
Recall that for $i\in[n]$, each observation $y_{i}$ is generated from model (1). Let ${\bm{\Psi}}\in[L]^{n}$ be the signal configuration vector, whose $i$-th entry stores the index of the signal underlying observation $y_{i}.$ That is, for $i\in[n]$ and $\ell\in[L]$, let $\Psi_{i}=\ell$ if and only if ${\bm{\beta}}^{(i)}$ equals the $\ell$th column of the signal matrix ${\bm{B}}$. We note that there is a one-to-one correspondence between ${\bm{\Psi}}$ and the change point vector ${\bm{\eta}}$. With a slight abuse of notation, we can rewrite the response vector $\bm{y}$ in a more general form:
@@
where $q$ is the known function in model (1), expanded to act row-wise on matrix inputs and to incorporate ${\bm{\Psi}}$, where $\left(q(\bm{X}\bm{B},{\bm{\Psi}},\bm{\varepsilon})\right)_{i}:=q(\bm{X}{\bm{\beta}}^{(\Psi_{i})},\varepsilon_{i})$ for $i\in[n]$. We note that the mixed linear regression model (Yi et al., 2014; Zhang et al., 2022; Tan and Venkataramanan, 2023) can also be written in the form in (5), with a crucial difference. In mixed regression, the components of ${\bm{\Psi}}$ are assumed to be drawn independently from some distribution on $[L]$, i.e., each $y_{i}$ is independently generated from one of the $L$ signals. In the change point setting, the entries of ${\bm{\Psi}}$ are dependent, since they change value only at entries $\eta_{1},\ldots,\eta_{(L^{*}-1)}$.
@@
$\begin{split}&\bm{\Theta}^{t}=\bm{X}\hat{\bm{B}}^{t}-\hat{{\bm{R}}}^{t-1}(\bm{F}^{t})^{\top}\,,\;\;\;\hat{{\bm{R}}}^{t}=g^{t}\left(\bm{\Theta}^{t},\bm{y}\right)\,,\\ &\bm{B}^{t+1}=\bm{X}^{\top}\hat{{\bm{R}}}^{t}-\hat{\bm{B}}^{t}(\bm{C}^{t})^{\top}\,,\;\;\;\hat{\bm{B}}^{t}=f^{t}\left(\bm{B}^{t}\right),\,\\ \end{split}$ (6)
@@
Crucially, the denoising functions $g^{t}$ and $f^{t}$ are not restricted to being separable. (Recall that a separable function acts row-wise identically on its input, i.e., $g:{\mathbb{R}}^{m\times L}\to{\mathbb{R}}^{m\times L}$ is separable if for all $\bm{U}\in{\mathbb{R}}^{m\times L}$ and $i\neq j$, we have $[g(\bm{U})]_{i}=g_{i}(\bm{U}_{i})=g_{j}(\bm{U}_{i})$.) Hence, in general, we have
@@
Our non-separable approach is required to handle the temporal dependence created by the change points. For example, if there was one change point uniformly in distributed in $[n]$, then $g^{t}(\bm{\Theta},\bm{y})$ should take into account that $y_{i}$ is more likely to have come from the the first signal for $i$ close to $1$, and from the second signal for $i$ close to $n$. This is in contrast to existing AMP algorithms for mixed regression (Tan and Venkataramanan, 2023), where under standard model assumptions, it suffices to consider separable denoisers.
@@
The memory terms $-\hat{{\bm{R}}}^{t-1}(\bm{F}^{t})^{\top}$ and $-\hat{\bm{B}}^{t}(\bm{C}^{t})^{\top}$ in our AMP algorithm (LABEL:eq:amp) debias the iterates $\bm{\Theta}^{t}$ and $\bm{B}^{t+1}$, and enable a succinct distributional characterization. In the high-dimensional limit as $n,p\to\infty$ (with $n/p\to\delta$), the empirical distributions of $\bm{\Theta}^{t}$ and $\bm{B}^{t+1}$ are quantified through the random variables $\bm{V}_{\bm{\Theta}}^{t}$ and $\bm{V}_{\bm{B}}^{t+1}$ respectively, where
@@
The matrices ${\bm{\rho}},{\bm{\nu}}_{\bm{\Theta}}^{t},{\bm{\nu}}_{\bm{B}}^{t+1}\in{\mathbb{R}}^{L\times L}$ are deterministic and defined below. The random matrices $\bm{Z}$, ${\bm{G}}_{\bm{\Theta}}^{t}$, and ${\bm{G}}_{\bm{B}}^{t+1}$ are independent of $\bm{X}$, and have i.i.d. rows following a Gaussian distribution. Namely, for $i\in[n]$ we have $wZ_{i}\mathbin{\overset{i.i.d}{\kern 0.0pt\sim}}{\sf N}({\bm{0}},{\bm{\rho}})$. For $i\in[n]$ and $s,r\geq 0$, $(\bm{G}_{\bm{\Theta}}^{t})_{i}\mathbin{\overset{i.i.d}{\kern 0.0pt\sim}}{\sf N}({\bm{0}},{\bm{\kappa}}_{\bm{\Theta}}^{t,t})$ with ${\rm{{Cov}}}((\bm{G}_{\bm{\Theta}}^{r})_{i},(\bm{G}_{\bm{\Theta}}^{s})_{i})={\bm{\kappa}}_{\bm{\Theta}}^{r,s}$. Similarly, for $j\in[p]$, $(\bm{G}_{\bm{B}}^{t})_{j}\mathbin{\overset{i.i.d}{\kern 0.0pt\sim}}{\sf N}({\bm{0}},{\bm{\kappa}}_{\bm{B}}^{t,t})$ with ${\rm{{Cov}}}((\bm{G}_{\bm{B}}^{r})_{j},(\bm{G}_{\bm{B}}^{s})_{j})={\bm{\kappa}}_{\bm{B}}^{r,s}$. The $L\times L$ deterministic matrices ${\bm{\nu}}_{\bm{\Theta}}^{t},{\bm{\kappa}}_{\bm{\Theta}}^{r,s},{\bm{\nu}}_{\bm{B}}^{t},$ and ${\bm{\kappa}}_{\bm{B}}^{r,s}$ are defined below via the state evolution recursion.
@@
The expectations above are taken with respect to $\bm{Z},\bm{V}_{\bm{\Theta}}^{t},\bm{V}_{\bm{\Theta}}^{s},\bm{V}_{\bm{B}}^{t+1}$ and $\bm{V}_{\bm{B}}^{s+1}$, and depend on $g^{t}$, $f^{t}$, $\bm{B}$, $\bm{\varepsilon}$, and ${\bm{\Psi}}$. The limits in (10)–(13) exist under suitable regularity conditions on $f_{t},g_{t}$, and on the limiting empirical distributions on $\bm{B},\bm{\varepsilon}$; see Appendix B. The dependence on $\bm{B}$, $\bm{\varepsilon}$ can also be removed under these conditions – this is discussed in the next subsection.

### v3

@@ Abstract
We consider the problem of localizing change points in a generalized linear model (GLM), a model that covers many widely studied problems in statistical learning including linear, logistic, and rectified linear regression. We propose a novel and computationally efficient approximate message passing (AMP) algorithm for estimating both the signals and the change point locations, and rigorously characterize its performance in the high-dimensional limit where the number of parameters $p$ is proportional to the number of samples $n$. This characterization is in terms of a state evolution recursion, which allows us to precisely compute performance measures such as the asymptotic Hausdorff error of our change point estimates, and allows us to tailor the algorithm to take advantage of any prior structural information of the signals and change points. Moreover, we show how our AMP iterates can be used to efficiently compute a Bayesian posterior distribution over the change point locations in the high-dimensional limit. We validate our theory via numerical experiments, and demonstrate the favorable performance of our estimators on both synthetic and real data in the settings of linear, logistic, and rectified linear regression.
Keywords: change point detection, high-dimensional regression, generalized linear models, multi-index models, approximate message passing, data segmentation
@@ 1 Introduction
Heterogeneity is a common feature of large, high-dimensional data sets. When the data are ordered by time, a simple form of heterogeneity is a change in the data generating mechanism at certain unknown instants of time. If these ‘change points’ were known, or estimated accurately, the data set could be partitioned into homogeneous subsets, each amenable to analysis via standard statistical techniques (Fryzlewicz, 2014). Models with change points have been studied in a variety of statistical contexts, such as the detection of changes in: signal means (Wang et al., 2020; Liu et al., 2021; Li et al., 2024); covariance structures (Cho and Fryzlewicz, 2014; Wang et al., 2021b); graphs (Londschien et al., 2023; Bhattacharjee et al., 2020; Fan and Guan, 2018); dynamic networks (Wang et al., 2021a); and functionals (Madrid Padilla et al., 2022). Change point models have found application in a range of fields including genomics (Braun et al., 2000), neuroscience (Aston and Kirch, 2012), causality (He et al., 2022), and economics (Andreou and Ghysels, 2002).
@@
Here, ${\bm{\beta}}^{(i)}\in{\mathbb{R}}^{p}$ is the unknown regression vector (signal) for the $i$th sample, $\bm{X}_{i}\in{\mathbb{R}}^{p}$ is the (known) covariate vector, $\varepsilon_{i}$ is noise, and $q:{\mathbb{R}}\times{\mathbb{R}}\to{\mathbb{R}}$ is a known function. We denote the unknown change points, that is, the sample indices where the regression vector changes, by $\eta_{1},\ldots,\eta_{L^{*}-1}$. Specifically, we have
$1=\eta_{0}<\eta_{1}<\dots<\eta_{L^{*}}=n+1,$
with ${\bm{\beta}}^{(i)}\neq{\bm{\beta}}^{(i-1)}$ if and only if $i\in\{\eta_{\ell}\}_{\ell=1}^{L^{*}-1}$. We note that $L^{*}$ is the number of distinct signals in the sequence $\{{\bm{\beta}}^{(i)}\}_{i=1}^{n}$, and $(L^{*}-1)$ is the number of change points. The number of change points is not known, but an upper bound $L$ on the value of $L^{*}$ is available. The goal is to estimate the change point locations as well as the $L^{*}$ signals. We would also like to quantify the uncertainty in these estimates, for example, via confidence sets or a posterior distribution. As we describe below, the model (1) covers many widely studied regression models including linear, logistic, and rectified linear regression.
@@ 1.1 Linear Regression with Change Points
Linear regression with change points in the high-dimensional regime (where the dimension $p$ is comparable to, or exceeds, the number of samples $n$) has been studied in a number of recent works, such as Lee et al. (2015); Leonardi and Bühlmann (2016); Kaul et al. (2019); Rinaldo et al. (2021); Xu et al. (2024); Li et al. (2023); Bai and Safikhani (2024). Most of these papers consider the setting where the signals are sparse (the number of non-zero entries in ${\bm{\beta}}^{(i)}\in{\mathbb{R}}^{p}$ is $o(p)$), and analyze procedures that combine the LASSO estimator (or a variant) with a partitioning technique such as dynamic programming. The work of Gao and Wang (2022) assumes sparsity on the difference between signals across a change point, and Cho et al. (2025); Liu et al. (2026) consider general non-sparse signals. The recent works of Liu et al. (2026); Yang et al. (2025) propose change point detection methods for high-dimensional linear models that do not rely on strict sub-Gaussian error assumptions, with performance comparable to that of Xu et al. (2024) in the case of light-tailed additive noise distributions.
@@ 1.2 Logistic Regression with Change Points
Logistic models are widely used for classification, and logistic regression with change points has been studied in number of contexts. In epidemiology, for example, logistic regression with unknown change points has been used to model the relationship between the continuous exposure variable and disease risk (Pastor‐Barriuso et al., 2003). In medicine, it has been used to identify relevant immune response biomarkers in patients with potentially infectious diseases (Fong et al., 2015).
@@ 1.3 Rectified Linear Regression with Change Points
Rectified linear regression is an important primitive in the theory of deep learning, and the problem without change points has been extensively studied in machine learning in recent years, both from the perspective of designing new algorithms and from that of computational hardness, as seen in the works of Soltanolkotabi (2017); Diakonikolas et al. (2021).
@@
Although existing procedures for high-dimensional change point regression such as Wang et al. (2023) incorporate sparsity-based constraints, they cannot be easily adapted to take advantage of other kinds of signal priors. Moreover, they are not well-equipped to exploit prior information on the change point locations. Bayesian approaches to change point detection have been studied in several works, such as those of Fearnhead (2006); Lungu et al. (2022), however they mainly focus on (low-dimensional) time-series.
@@
Although our theoretical results do not explicitly need assumptions on the separation between adjacent change points, they are most interesting when the separation is of order $n$, that is, $\Delta:=\min_{\ell\in[L]}(\eta_{\ell}-\eta_{\ell-1})=O(n)$. This separation is natural in our regime, where the number of samples $n$ is proportional to $p$ and the number of degrees of freedom in the signals also grows linearly in $p$. Existing results on high-dimensional change point regression usually assume signals that are $s$-sparse with $s$ sufficiently small, and demonstrate change point estimators that are consistent when $\Delta=\omega\big{(}s\log p/\kappa^{2}\big{)}$, where $\kappa$ is a constant determined by the separation between the signals (Wang et al., 2021c; Li et al., 2023; Wang et al., 2023). In contrast, we do not assume signal sparsity that is sublinear in $n$, so the change point estimation error will not converge to zero unless $n/p\to\infty$. We therefore quantify the AMP performance via precise asymptotics for the estimation error and the limiting posterior distribution.
@@
AMP, a family of iterative algorithms first proposed for linear regression (Kabashima, 2003; Donoho et al., 2009; Krzakala et al., 2012), has been applied to a variety of high-dimensional estimation problems including estimation in generalized linear models (Rangan, 2011; Schniter and Rangan, 2014; Maillard et al., 2020; Mondelli and Venkataramanan, 2021) and low-rank matrix estimation (Fletcher and Rangan, 2018; Lesieur et al., 2017; Montanari and Venkataramanan, 2021; Barbier et al., 2020). An attractive feature of AMP algorithms is that under suitable model assumptions, their performance in the high-dimensional limit can be characterized by a succinct deterministic recursion called state evolution. The state evolution characterization has been used to show that AMP achieves Bayes-optimal performance for some models (Deshpande and Montanari, 2014; Donoho et al., 2013; Barbier et al., 2019). We refer the reader to Appendix G for further background on AMP for Generalized Linear Models.
An important feature of our AMP algorithm, in contrast to the above works, is the use of non-separable denoising functions. (We say a function $g:{\mathbb{R}}^{m\times L}\to{\mathbb{R}}^{m\times L}$ is separable if it acts row-wise identically on the input matrix.) Even with simplifying assumptions on the signal and noise distributions, non-separable AMP denoisers are required to handle the heterogeneous dependence structure caused by change points, and allow for precise uncertainty quantification around possible change point locations. Our main state evolution result (Theorem 1) leverages recent results by Berthier et al. (2019) and Gerbelot and Berthier (2023) for AMP with non-separable denoisers. Non-separable denoisers for AMP have been been studied for generalized linear models and for matrix tensor product models, to exploit the dependence within the signal (Som and Schniter, 2012; Metzler et al., 2016; Ma et al., 2019; Rossetti and Reeves, 2023) or between the covariates (Zhang et al., 2024). In our setting, non-separable denoisers are required to exploit the heterogeneous dependence structure caused by change points.
@@ 2 Preliminaries
We let $n]=\{1,2,\dots,n\}$. We use boldface notation for matrices and vectors. For vectors $\bm{x},\bm{y}\in{\mathbb{R}}^{n}$, we write $\bm{x}\leq\bm{y}$ if $x_{i}\leq y_{i}$ for all $i\in[n]$, and let $\bm{x},\bm{y}]:=\{\bm{v}\in{\mathbb{R}}^{n}:x_{i}\leq v_{i}\leq y_{i}\;\forall i\in[n]\}$. For a matrix $\bm{A}\in{\mathbb{R}}^{m\times L}$ and $i\in[m],j\in[L]$, we let $\bm{A}_{[i,:]}$ and $\bm{A}_{[:,j]}$ denote its $i$th row and $j$th column respectively. Similarly, for a vector ${\bm{\psi}}\in[L]^{n}$, we let $\bm{A}_{[:,{\bm{\psi}}]}\in{\mathbb{R}}^{m}$ denote the vector whose $i$-th entry is $\bm{A}_{i,\psi_{i}}$. For two sequences (in $n$) of random variables $X_{n}$, $Y_{n}$, we write $X_{n}\stackrel{{\scriptstyle{\mathbb{P}}}}{{\simeq}}Y_{n}$ when their difference converges in probability to 0, that is, $\lim_{n\to\infty}{\mathbb{P}}(|X_{n}-Y_{n}|>\epsilon)=0$ for any $\epsilon>0$. Denote the covariance matrix of random vector $\bm{Z}\in\mathbb{R}^{q}$ as ${\rm{{Cov}}}(\bm{Z})\in\mathbb{R}^{q\times q}$. We refer to all random elements, including vectors and matrices, as random variables. When referring to probability densities, we include probability mass functions, with integrals interpreted as sums when the distribution is discrete. Given a matrix $\bm{B}\in{\mathbb{R}}^{p\times L}$, the empirical distribution of the entries of $\bm{B}$ is the measure $\mu_{p}(A):=\frac{1}{p}|\{\bm{B}_{j}\in A:j\in[p]\}|$ for any measurable set $A\subseteq{\mathbb{R}}^{L}$. We say that the empirical distribution of the entries of $\bm{B}$ converges weakly to the law ${\mathbb{P}}_{\bar{\bm{B}}}$ if $\frac{1}{p}\sum_{j=1}^{p}f(\bm{B}_{j})\to{\mathop{\mathbb{E}}}_{{\mathbb{P}}_{\bar{\bm{B}}}}[f(\bar{\bm{B}})]$ as $p\to\infty$ for all bounded continuous functions $f:{\mathbb{R}}^{L}\to{\mathbb{R}}$.
@@
We will state our results in terms of uniformly pseudo-Lipschitz functions (Berthier et al., 2019). For $C>0$ and $r\in[1,\infty)$, let $PL_{n,m,q}(r,C)$ be the set of functions $\phi:{\mathbb{R}}^{n\times q}\to{\mathbb{R}}^{m\times q}$ such that
@@ 3 AMP Algorithm and Main Results
[…]
@@
where $q$ is the known function in model (1), expanded to act row-wise on matrix inputs and to incorporate $\bm{\Psi}$}, where $\left(q(\bm{X}\bm{B},{\bm{\Psi}},\bm{\varepsilon})\right)_{i}:=q(\bm{X}{\bm{\beta}}^{(\Psi_{i})},\varepsilon_{i})$ for $i\in[n]$. We note that the mixed linear regression model (Yi et al., 2014; Zhang et al., 2022; Tan and Venkataramanan, 2023) can also be written in the form in (5), with a crucial difference. In mixed regression, the components of $\bm{\Psi}$} are assumed to be drawn independently from some distribution on $[L]$, that is, each $y_{i}$ is independently generated from one of the $L$ signals. In the change point setting, the entries of $\bm{\Psi}$} are dependent, since they change value only at indices $\eta_{1},\ldots,\eta_{(L^{*}-1)}$.
@@
[…]
@@
Crucially, the denoising functions $g^{t}$ and $f^{t}$ are not restricted to being separable. (Recall that a separable function acts row-wise identically on its input, that is, $g:{\mathbb{R}}^{m\times L}\to{\mathbb{R}}^{m\times L}$ is separable if for all $\bm{U}\in{\mathbb{R}}^{m\times L}$ and $i\neq j$, we have $g(\bm{U})]_{i}=g_{i}(\bm{U}_{i})=g_{j}(\bm{U}_{i})$.) Hence, in general, we have
@@
Our non-separable approach is required to handle the heterogeneous dependence structure created by the change points. For example, if there was one change point uniformly distributed in $[n]$, then $g^{t}(\bm{\Theta},\bm{y})$ should take into account that $y_{i}$ is more likely to have come from the the first signal for $i$ close to $1$, and from the second signal for $i$ close to $n$. This is in contrast to existing AMP algorithms for mixed regression (Tan and Venkataramanan, 2023), where under standard model assumptions, it suffices to consider separable denoisers.
@@
The memory terms $\hat{{\bm{R}}}^{t-1}(\bm{F}^{t})^{\top}$ and $\hat{\bm{B}}^{t}(\bm{C}^{t})^{\top}$ in our AMP algorithm (LABEL:eq:amp) debias the iterates $\bm{\Theta}^{t}$ and $\bm{B}^{t+1}$, and enable a succinct distributional characterization. In the high-dimensional limit as $n,p\to\infty$ (with $n/p\to\delta$), the empirical distributions of $\bm{\Theta}^{t}$ and $\bm{B}^{t+1}$ are quantified through the random variables $\bm{V}_{\bm{\Theta}}^{t}$ and $\bm{V}_{\bm{B}}^{t+1}$ respectively, where
@@
The matrices ${\bm{\rho}},{\bm{\nu}}_{\bm{\Theta}}^{t},{\bm{\nu}}_{\bm{B}}^{t+1}\in{\mathbb{R}}^{L\times L}$ are deterministic and defined below. The random matrices $\bm{Z}$, ${\bm{G}}_{\bm{\Theta}}^{t}$, and ${\bm{G}}_{\bm{B}}^{t+1}$ are independent of $\bm{X}$, and have i.i.d. rows following a Gaussian distribution. Namely, for $i\in[n]$ we have $Z_{i}\mathbin{\overset{i.i.d}{\kern 0.0pt\sim}}{\sf N}({\bm{0}},{\bm{\rho}})$. For $i\in[n]$ and $s,r\geq 0$, $(\bm{G}_{\bm{\Theta}}^{t})_{i}\mathbin{\overset{i.i.d}{\kern 0.0pt\sim}}{\sf N}({\bm{0}},{\bm{\kappa}}_{\bm{\Theta}}^{t,t})$ with ${\rm{{Cov}}}((\bm{G}_{\bm{\Theta}}^{r})_{i},(\bm{G}_{\bm{\Theta}}^{s})_{i})={\bm{\kappa}}_{\bm{\Theta}}^{r,s}$. Similarly, for $j\in[p]$, $(\bm{G}_{\bm{B}}^{t})_{j}\mathbin{\overset{i.i.d}{\kern 0.0pt\sim}}{\sf N}({\bm{0}},{\bm{\kappa}}_{\bm{B}}^{t,t})$ with ${\rm{{Cov}}}((\bm{G}_{\bm{B}}^{r})_{j},(\bm{G}_{\bm{B}}^{s})_{j})={\bm{\kappa}}_{\bm{B}}^{r,s}$. The $L\times L$ deterministic matrices ${\bm{\nu}}_{\bm{\Theta}}^{t},{\bm{\kappa}}_{\bm{\Theta}}^{r,s},{\bm{\nu}}_{\bm{B}}^{t},$ and ${\bm{\kappa}}_{\bm{B}}^{r,s}$ are defined below via the state evolution recursion.
@@
The expectations above are taken with respect to $\bm{Z},\bm{V}_{\bm{\Theta}}^{t},\bm{V}_{\bm{\Theta}}^{s},\bm{V}_{\bm{B}}^{t+1}$ and $\bm{V}_{\bm{B}}^{s+1}$, and depend on $g^{t}$, $f^{t}$, $\bm{B}$, $\bm{\varepsilon}$, and $\bm{\Psi}$}. The limits in (10)–(13) exist under suitable regularity conditions on $f_{t},g_{t}$, and on the limiting empirical distributions on $\bm{B},\bm{\varepsilon}$; see Appendix B. The parametric dependence on $\bm{B}$, $\bm{\varepsilon}$ can also be removed under these conditions—this is discussed in the next subsection.